Soybean meal is 16% protein; cornmeal is 8% protein. How many pounds of each should be mixed together in order to get 320lb mixture that is 10% protein?
Answer by checkley77(12569)

.16s+.08(320-s)=.10*320
.16s+25.6-.08s=32
.08s=32-25.6
.08s=6.4
s=6.4/.08
s=80 pounds of soybeans are used.
320-80=240 pounds of cornmeal are used.
Proof:
.16*80+.08*240=.10*320
12.8+19.2=32
32=32